OpenStudy (anonymous):

percent increase or decrease : before 40 after 65 what are the steps to do this i forgot???

OpenStudy (karatechopper):

amount of change/ original

OpenStudy (karatechopper):

so it would be 65-40/40

OpenStudy (anonymous):

so its 40/65??

OpenStudy (karatechopper):

15/40

OpenStudy (karatechopper):

simplify that..3/8

OpenStudy (karatechopper):

use a calculator...

OpenStudy (karatechopper):

3/8=0.375 move decimal 2 steps to right 37.5% increse

OpenStudy (anonymous):

8/13????

OpenStudy (karatechopper):

nooo

OpenStudy (karatechopper):

lets take this step by step k?

OpenStudy (anonymous):

ok

OpenStudy (karatechopper):

so the formula is change in amount/ orginal

OpenStudy (karatechopper):

so the change in amount is the change between 65 and 40 so 65-40/original

OpenStudy (karatechopper):

and the original is 40, because u said before.

OpenStudy (karatechopper):

so now the equation is 65-40/40

OpenStudy (karatechopper):

so if u simplify that... it is 15/40

OpenStudy (karatechopper):

and if i simplify that to work with better, it becomes 3/8

OpenStudy (karatechopper):

then i use the calculator asking what 3/8 is..

OpenStudy (karatechopper):

it says 3/8=0.375

OpenStudy (karatechopper):

and since u are looking for percent i move the decimal 2 steps to the right.

OpenStudy (karatechopper):

so not 3/8= 37.5%

OpenStudy (karatechopper):

and if u are counting from 40 to 65, u are INCREASING your count...

OpenStudy (karatechopper):

therefore it is a 37.5% increase

OpenStudy (karatechopper):

get it?

OpenStudy (anonymous):

im stuck at where i subtract then what do i do?

OpenStudy (anonymous):

Percentage Increase or Decrease \[=\frac{amount.of.change}{original.price} \times100\]\[=\frac{NP-OP}{OP} \times 100\] where NP is your new price and OP is the old price

OpenStudy (anonymous):

so in your question what was the old price and what was the new price?

OpenStudy (anonymous):

40 old price new price 65

OpenStudy (anonymous):

okay good so OP=40 and NP=65. try and put these values into the equation i gave above and what answer do you get?

OpenStudy (anonymous):

so i subtract 65 -40 and i get 25 then what do i do multiply with 100?

OpenStudy (anonymous):

well let's multiply by 100 last because all that does is change our answer to percentage. so you have 25 now you need to divide by the original price

OpenStudy (anonymous):

0.625 i got

OpenStudy (anonymous):

good now the last step is to times by 100

OpenStudy (anonymous):

62.5

OpenStudy (anonymous):

perfect so your answer is 62.5%

OpenStudy (anonymous):

so how do i know if its increasing or decreasing?

OpenStudy (anonymous):

if the percentage is positive it's increasing if it's negative it's decreasing

OpenStudy (anonymous):

thanks!!!!!!!!!! ALOT!

OpenStudy (anonymous):

no worries (: